The gross payoffs (X) from the investment have the following probability mass function.

x p(x)
0 0.5
2000 0.2
5000 0.3

The expected gross payoff (mean payoff) is:...............
Answer #1

Solution :

given table

x p(x)
0 0.5
2000 0.2
5000 0.3

expected value = E(x)= X * P(X)

= 0 * 0.5+ 2000 * 0.2+ 5000 *0.3  

= ( 0+ 400+ 1500 )

= 1900

The expected gross payoff (mean payoff) is:......1900